# Almost sure convergence rates of stochastic gradient methods under gradient domination

### v1

@@ Appendix F Proof of Section 5
$\begin{split}\tilde{D}_{n+1}&\leq D_{n+1}\mathbf{1}_{\Omega_{n}}\\ &\leq D_{n}\mathbf{1}_{\Omega_{n}}-\gamma_{n}\mathbf{1}_{\Omega_{n}}\lVert\nabla f(X_{n})\rVert^{2}+\gamma_{n}\mathbf{1}_{\Omega_{n}}\xi_{n+1}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2}\\ &\leq D_{n}\mathbf{1}_{\Omega_{n}}-\gamma_{n}c^{2}(f(X_{n})-l)^{2\beta}\mathbf{1}_{\Omega_{n}}+\gamma_{n}\mathbf{1}_{\Omega_{n}}\xi_{n+1}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2}\\ &=\tilde{D}_{n}-\gamma_{n}c^{2}\tilde{D}_{n}^{2\beta}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2},\\ &=(1-\gamma_{n}^{q}c^{2})\tilde{D}_{n}+\gamma_{n}c^{2}(\gamma_{n}^{1-q}\tilde{D}_{n}-\tilde{D}_{n}^{2\beta})+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}\mathbf{1}_{\Omega_{n}}\\ &\leq(1-\gamma_{n}^{q}c^{2})\tilde{D}_{n}+(2\beta)^{-\frac{1}{2\beta-1}}(1-\frac{1}{2\beta}){c^{2}}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}\mathbf{1}_{\Omega_{n}}\\ &\leq\tilde{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}^{q}c^{2})+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{k}}+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{k}},\end{split}$ (35)
for $\tilde{c}=(2\beta)^{-\frac{1}{2\beta-1}}(1-\frac{1}{2\beta})c^{2}$ from the function trick (12) which we applied in the forth inequality. We also used that the unified gradient domination property holds for all $X_{k}$, $k\leq n$ on the event $\Omega_{n}$ ∎
@@
$M_{n}=\sum_{k=1}^{n}\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{k}}\quad\textrm{ and }\quad S_{n}=\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{k}}.$
Then, $(M_{n})$ is a $(\mathcal{F}_{n-1})$-martingale with zero mean and $(S_{n})$ is a $(\mathcal{F}_{n-1})$-sub-martingale by Assumption 2.4. Note that by the choice of $\gamma_{n}$ we have that $\sum_{n}\gamma_{n}^{2}<\infty$ and hence $\mathbb{E}[S_{n}]<\infty$ for all $n\in\mathbb{N}$.
Next, define $R_{n}=M_{n}^{2}+S_{n}$ for every $n\in\mathbb{N}$. Moreover, let
@@
which is an $\mathcal{F}_{n}$-measurable event on $(\Omega,\mathcal{F},\mathbb{P})$. We define $R_{0}=0$ such that $E_{0}=\Omega$.
Now let $\hat{E}_{n}=E_{n}\cap C_{n}$, then we will first show, that $\hat{E}_{n}$ fulfills the property $\hat{E}_{n}\subset\Omega_{n+1}$ for all $n\in\mathbb{N}$ in Lemma F.6 and then that $\hat{E}_{n}$ occurs with probability at least $1-\delta$ in Lemma F.8.
@@
By triangle inequality we have that $\inf_{x^{\ast}\in\mathcal{X}^{\ast}}||y-x^{\ast}||\leq\frac{3\mathbf{r}}{4}$, i.e there exists $x^{\ast}\in\mathcal{X}^{\ast}$ such that $||y-x^{\ast}||\leq\frac{3\mathbf{r}}{4}$. Suppose now, that $\inf_{x^{\ast}\in\mathcal{X}^{\ast}}||y-x^{\ast}||\geq\frac{\mathbf{r}}{2}$, this means that $y\in\bigcup_{x^{\ast}\in\mathcal{X}^{\ast}}\mathcal{B}_{\frac{3r}{4}}(x^{\ast})\setminus\bigcup_{x^{\ast}\in\mathcal{X}^{\ast}}\tilde{\mathcal{B}}_{\frac{\mathbf{r}}{2}}(x^{\ast})$. By the definition of $s=\inf\left\{f(z)-l\,:\;z\in\bigcup_{x^{\ast}\in\mathcal{X}^{\ast}}\mathcal{B}_{\frac{3\mathbf{r}}{4}}(x^{\ast})\setminus\bigcup_{x^{\ast}\in\mathcal{X}^{\ast}}\tilde{\mathcal{B}}_{\frac{\mathbf{r}}{2}}(x^{\ast})\right\}$ this contradicts the second assumption $f(y)-l<s$. ∎
@@
almost surely. We prove the assertion by induction. Let $\omega\in\hat{E}_{n}$. Since $\hat{E}_{n}\subset\hat{E}_{n-1}\subset\Omega_{n}$ by induction assumption, we have $\omega\in\Omega_{n}$ and thus $\omega\in\Omega_{k}$ for all $k\leq n$. We will apply Lemma F.5 with $x=X_{n}(\omega)$ and $y=X_{n+1}(\omega)$. By definition it holds that $\omega\in\hat{E}_{n}$ implies condition $3$ and $\omega\in\Omega_{n}$ implies condition $1$ of Lemma F.5. It remains to show condition $2$, then it follows that $\inf_{x^{\ast}\in\mathcal{X}^{\ast}}||X_{n+1}(\omega)-x^{\ast}||<\frac{\mathbf{r}}{2}$, i.e. $X_{n+1}(\omega)\in\mathcal{U}$ and by $\omega\in\Omega_{n}$ we deduce $\omega\in\Omega_{n+1}$.
@@
$\beta=\frac{1}{2}$: The inequality (30) and the induction hypothesis yield
@@
$\leq D_{1}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}c)+\sum_{k=1}^{n}\gamma_{k}\xi_{k+1}(\omega)\mathbf{1}_{\Omega_{k}}(\omega)+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}\mathbf{1}_{\Omega_{k}}(\omega)$
@@
$\leq 2\epsilon+\sqrt{\epsilon}<s.$
$\beta\in(\frac{1}{2},1]$: Similarly, we obtain from (45)
@@
$\leq D_{1}(\omega)\mathbf{1}_{\Omega_{1}}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}^{q}c)+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\gamma_{k}\xi_{k+1}(\omega)+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}$
@@
We used in both cases that that $\prod_{k=1}^{n}(1-\gamma_{k}^{(q)}c)\leq 1$ and the choice of $\epsilon$ such that $2\epsilon+\sqrt{\epsilon}<s$. This proves that condition $2$ in Lemma F.5 is also satisfied which concludes the induction.
c) By definition it holds that $E_{n}=E_{n-1}\setminus(E_{n-1}\setminus E_{n})=E_{n-1}\setminus\tilde{E}_{n}$. Then we have
@@
$=\gamma_{n}^{2}\xi_{n+1}^{2}\mathbf{1}_{\Omega_{n}}+2\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}M_{n-1}^{2}+\gamma_{n}^{2}\frac{L}{2}\lVert V_{n+1}(X_{n})\rVert^{2}\mathbf{1}_{\Omega_{n}}.$
@@
$\mathbb{E}[\xi_{n+1}M_{n-1}^{2}\mathbf{1}_{\Omega_{n}}]$ $=\mathbb{E}[\mathbb{E}[\xi_{n+1}|\mathcal{F}_{n}]M_{n-1}^{2}\mathbf{1}_{\Omega_{n}}]=0,$
@@
Putting all together we obtain the claim
@@
Lemma F.7.
Let $\delta>0$ be a tolerance level and $\gamma_{n}\leq\gamma_{1}$ be sufficiently small such that $\sum_{n=1}^{\infty}\gamma_{n}^{2}<\frac{\delta\epsilon}{2(G^{2}C^{2}+G^{2}+C)}$ and the condition in Lemma F.6 is fulfilled. Then, we have
@@
On the other hand it follows from Lemma F.10 that
@@
Lemma F.8.
Let $\delta>0$ be a tolerance level and $\gamma_{n}\leq\gamma_{1}$ be sufficiently small such that the condition in Lemma F.6 and Lemma F.7 are fulfilled. Moreover, we suppose $\gamma_{1}$ small enough such that $\frac{4C}{\mathbf{r}^{2}}\sum_{k=1}^{n}\gamma_{k}^{2}\leq\frac{\delta}{2}$. Then, we have
@@
By Lemma F.8, we have $\mathbb{P}(E_{n})\geq 1-\frac{\delta}{2}$. Moreover, by the additional step size assumption and Markov’s inequality we deduce that
@@
Hence, using Lemma F.8 we obtain
@@
(ii): We prove that $\tilde{D}_{n}=D_{n}\mathbf{1}_{\Omega_{n}}\in o(1/n^{1-\eta})$, then the claim follows since $\mathbf{1}_{\Omega_{\mathcal{U}}}\leq\mathbf{1}_{\Omega_{n}}$ almost surely.
From the proof of Lemma F.4 we have
$\tilde{D}_{n+1}\leq\tilde{D}_{n}-\gamma_{n}c\tilde{D}_{n}^{2\beta}+\gamma_{n}\xi_{n}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n}\rVert^{2}\mathbf{1}_{\Omega_{n}}.$
@@
$\leq\tilde{D}_{n}-\gamma_{n}c\tilde{D}_{n}^{2\beta}+LC\gamma_{n}^{2},$
where we have used that $D_{n}$ and $\mathbf{1}_{\Omega_{n}}$ are $\mathcal{F}_{n}$-measurable and $E[\lVert V_{n+1}(X_{n})\rVert^{2}|\mathcal{F}_{n}]\leq C$ from (ABC) with $A=B=0$. By our step size choice we can apply Lemma 3.1 to obtain Claim (ii).
@@
$\beta=\frac{1}{2}$: We have from Lemma F.4 Equation 45 that
$\tilde{D}_{n+1}$ $\leq(1-\gamma_{n}c^{2})\tilde{D}_{n}+\gamma_{n}\xi_{n}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}.$
@@
where we used (37) in the first inequality. By our choice of $\gamma_{n}$ there exists $\tilde{c}>0$ and $N>0$ such that $\gamma_{n}c^{2}-\frac{1-\eta}{n}+\frac{(1-\eta)\gamma_{n}c^{2}}{n}\geq\tilde{c}\gamma_{n}$ for all $n\geq N$. Thus, for all $n\geq N$
@@
by (16) in Lemma 3.1 Hence, we apply Lemma A.3 to prove that $\lim_{n\to\infty}w_{n}=0$. By the definition of $w_{n}$ we have verified that $\mathbb{E}[(f(X_{n})-l)\mathbf{1}_{\Omega_{\mathcal{U}}}]\leq\mathbb{E}[\tilde{D}_{n}]\in o(\frac{1}{n^{1-\eta}})$
$\beta\in(\frac{1}{2},1]$: From Lemma F.4 Equation 45 we have
$\tilde{D}_{n+1}$ $\leq(1-\gamma_{n}^{q}c^{2})\tilde{D}_{n}+\tilde{c}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+\gamma_{n}\xi_{n}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2},$
for $\tilde{c}=(2\beta)^{-\frac{1}{2\beta-1}}(1-\frac{1}{2\beta})c^{2}$. Next we multiply with $(n+1)^{1-\eta}$ and use (9) to obtain
@@
$\leq(n+1)^{1-\eta}(1-\gamma_{n}^{q}c^{2})\mathbb{E}[\tilde{D}_{n}]+(n+1)^{1-\eta}\tilde{c}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+(n+1)^{1-\eta}LC\gamma_{n}^{2}$
@@
because we choose the auxiliary parameter $q$ as in the proof of Lemma 3.1 where we showed in (16) and (17) that
@@
The proof will be similar to the previous section. Instead of assuring that $(X_{n})$ remains close to the set where we could guarantee the unified gradient domination property, it is now sufficient that $f(X_{n})$ remains close to $f^{\ast}$ by the different definition of gradient domination definition in $f^{\ast}$. This will simplify the proof. Moreover, we may again assume w.l.o.g. the uniform second moment bounds, i.e. $A=B=0$, instead of the more general (ABC) condition by the same argument as above but on the level sets.
@@
In the following, we fix the notation $D_{n}:=f(X_{n})-f^{\ast}$ and $\tilde{D}_{n}:=D_{n}\mathbf{1}_{\Omega_{n}}$ and obtain the parallel result to Lemma F.4.
Lemma F.9.

### v3

@@ Appendix D Proofs of Section 5
$\begin{split}D_{n+1}\mathbf{1}_{\Omega_{n}}&\leq D_{n}\mathbf{1}_{\Omega_{n}}-\gamma_{n}\mathbf{1}_{\Omega_{n}}\lVert\nabla f(X_{n})\rVert^{2}+\gamma_{n}\mathbf{1}_{\Omega_{n}}\xi_{n+1}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2}\\ &\leq D_{n}\mathbf{1}_{\Omega_{n}}-\gamma_{n}c^{2}(f(X_{n})-l)^{2\beta}\mathbf{1}_{\Omega_{n}}+\gamma_{n}\mathbf{1}_{\Omega_{n}}\xi_{n+1}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2}\\ &=D_{n}\mathbf{1}_{\Omega_{n}}-\gamma_{n}c^{2}D_{n}^{2\beta}\mathbf{1}_{\Omega_{n}}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\mathbf{1}_{\Omega_{n}}\lVert V_{n+1}(X_{n})\rVert^{2},\\ &=(1-\gamma_{n}^{q}c^{2}){D}_{n}\mathbf{1}_{\Omega_{n}}+\gamma_{n}c^{2}(\gamma_{n}^{1-q}{D}_{n}-{D}_{n}^{2\beta})\mathbf{1}_{\Omega_{n}}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}\mathbf{1}_{\Omega_{n}}\\ &\leq(1-\gamma_{n}^{q}c^{2}){D}_{n}\mathbf{1}_{\Omega_{n}}+(2\beta)^{-\frac{1}{2\beta-1}}(1-\frac{1}{2\beta}){c^{2}}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}\mathbf{1}_{\Omega_{n}}\\ &\leq{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}^{q}c^{2})\mathbf{1}_{\Omega_{n}}+\tilde{c}\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}\\ &\quad+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{n}}+\frac{L}{2}\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{n}}\\ &\leq{D}_{1}\prod_{k=1}^{n}(1-\gamma_{k}^{q}c^{2})\mathbf{1}_{\Omega_{n}}+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{n}}\\ &\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k})\rVert^{2}\mathbf{1}_{\Omega_{n}},\end{split}$ (33)
for $\tilde{c}=(2\beta)^{-\frac{1}{2\beta-1}}(1-\frac{1}{2\beta})c^{2}$ from the function trick Equation 20 which we applied in the forth inequality. We also used that the unified gradient domination property holds for all $X_{k}$, $k\leq n$ on the event $\Omega_{n}$ ∎
@@
$M_{n}=\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{k}},\quad M_{n}^{(q)}=\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}\xi_{k+1}\mathbf{1}_{\Omega_{k}}$
[…]
Then, $(M_{n})$ and $(M_{n}^{(q)})$ are $(\mathcal{F}_{n+1})$-martingales with zero mean and $(S_{n})$ is a $(\mathcal{F}_{n+1})$-sub-martingale by Assumption 2.4. Note that by the choice of $\gamma_{n}$ we have that $\sum_{n}\gamma_{n}^{2}<\infty$ and hence $\mathbb{E}[S_{n}]<\infty$ for all $n\in\mathbb{N}$.
Next, define $R_{n}=M_{n}^{2}+S_{n}$ and $R_{n}=(M_{n}^{(q)})^{2}+S_{n}$ respectively (with some abuse of notation), for every $n\in\mathbb{N}$. Moreover, let
@@
which is an $\mathcal{F}_{n+1}$-measurable event on $(\Omega,\mathcal{F},\mathbb{P})$. We define $R_{0}=0$ such that $E_{0}=\Omega$.
Now let $\hat{E}_{n}=E_{n}\cap C_{n}$, then we will first show, that $\hat{E}_{n}$ fulfills the property $\hat{E}_{n}\subset\Omega_{n+1}$ for all $n\in\mathbb{N}$ in Lemma D.6 and then that $\hat{E}_{n}$ occurs with probability at least $1-\delta$ in Lemma D.8.
@@
By triangle inequality we have that $\inf_{x^{\ast}\in\mathcal{X}^{\ast}}||y-x^{\ast}||\leq\frac{3\mathbf{r}}{4}$, i.e there exists $x^{\ast}\in\mathcal{X}^{\ast}$ such that $||y-x^{\ast}||\leq\frac{3\mathbf{r}}{4}$. Suppose now, that $\inf_{x^{\ast}\in\mathcal{X}^{\ast}}||y-x^{\ast}||\geq\frac{\mathbf{r}}{2}$, this means that $y\in\bigcup_{x^{\ast}\in\mathcal{X}^{\ast}}\mathcal{B}_{\frac{3r}{4}}(x^{\ast})\setminus\bigcup_{x^{\ast}\in\mathcal{X}^{\ast}}\tilde{\mathcal{B}}_{\frac{\mathbf{r}}{2}}(x^{\ast})$. By the definition of
$s=\inf\left\{f(z)-l\,:\;z\in\bigcup_{x^{\ast}\in\mathcal{X}^{\ast}}\mathcal{B}_{\frac{3\mathbf{r}}{4}}(x^{\ast})\setminus\bigcup_{x^{\ast}\in\mathcal{X}^{\ast}}\tilde{\mathcal{B}}_{\frac{\mathbf{r}}{2}}(x^{\ast})\right\}$
this contradicts the second assumption $f(y)-l<s$. ∎
@@
almost surely. We prove the assertion by induction. Let $\omega\in\hat{E}_{n}$. Since $\hat{E}_{n}\subset\hat{E}_{n-1}\subset\Omega_{n}$ by induction assumption, we have $\omega\in\Omega_{n}$ and thus $\omega\in\Omega_{k}$ for all $k\leq n$. We will apply Lemma D.5 with $x=X_{n}(\omega)$ and $y=X_{n+1}(\omega)$. By definition it holds that $\omega\in\hat{E}_{n}$ implies condition $3$ and $\omega\in\Omega_{n}$ implies condition $1$ of Lemma D.5. It remains to show condition $2$, then it follows that $\inf_{x^{\ast}\in\mathcal{X}^{\ast}}||X_{n+1}(\omega)-x^{\ast}||<\frac{\mathbf{r}}{2}$, i.e. $X_{n+1}(\omega)\in\mathcal{U}$ and by $\omega\in\Omega_{n}$ we deduce $\omega\in\Omega_{n+1}$.
@@
$\beta=\frac{1}{2}$: The inequality Equation 30 and the induction hypothesis yield
@@
$\leq D_{1}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}c^{2})+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}(\omega)\mathbf{1}_{\Omega_{n}}(\omega)$
$\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}\mathbf{1}_{\Omega_{n}}(\omega)$
$=D_{1}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}c^{2})+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}(\omega)\mathbf{1}_{\Omega_{k}}(\omega)$
$\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}\mathbf{1}_{\Omega_{k}}(\omega)$
@@
$\leq 2\epsilon+\sqrt{\epsilon}<s,$
where the equation in the third line is due to $\omega\in\Omega_{k}$ for all $k\leq n$ by induction.
$\beta\in(\frac{1}{2},1]$: Similarly, we obtain from Equation 42
@@
$\leq D_{1}(\omega)\mathbf{1}_{\Omega_{n}}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}^{q}c^{2})+\tilde{c}\sum_{k=1}^{n}\gamma_{k}^{\frac{2\beta q-1}{2\beta-1}}+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}^{q}c^{2})\right)\gamma_{k}\xi_{k+1}(\omega)\mathbf{1}_{\Omega_{n}}(\omega)$
$\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}\mathbf{1}_{\Omega_{n}}(\omega)$
$=D_{1}(\omega)\prod_{k=1}^{n}(1-\gamma_{k}^{q}c)+\sum_{k=1}^{n}\left(\prod_{j=k}^{n}(1-\gamma_{j}c^{2})\right)\gamma_{k}\xi_{k+1}(\omega)\mathbf{1}_{\Omega_{k}}(\omega)$
$\quad+\frac{L}{2}\sum_{k=1}^{n}\gamma_{k}^{2}\lVert V_{k+1}(X_{k}(\omega))\rVert^{2}\mathbf{1}_{\Omega_{k}}(\omega)$
@@
We used in both cases that that $\prod_{k=1}^{n}(1-\gamma_{k}^{q}c)\leq 1$ and the choice of $\epsilon$ such that $2\epsilon+\sqrt{\epsilon}<s$. This proves that condition $2$ in Lemma D.5 is also satisfied which concludes the induction.
c) Without loss of generality we consider the case $\beta=1/2$. The computations for $\beta\in(1/2,1]$ follow in line by replacing $M_{n}$ with $M_{n}^{(q)}$. By definition it holds that $E_{n}=E_{n-1}\setminus(E_{n-1}\setminus E_{n})=E_{n-1}\setminus\tilde{E}_{n}$. Then we have
@@
$=\gamma_{n}^{2}(1-\gamma_{n}c^{2})^{2}\xi_{n+1}^{2}\mathbf{1}_{\Omega_{n}}+2\gamma_{n}(1-\gamma_{n}c)\xi_{n+1}\mathbf{1}_{\Omega_{n}}M_{n-1}+\gamma_{n}^{2}\frac{L}{2}\lVert V_{n+1}(X_{n})\rVert^{2}\mathbf{1}_{\Omega_{n}}.$
@@
$\mathbb{E}[\xi_{n+1}(1-\gamma_{n}c)M_{n-1}\mathbf{1}_{\Omega_{n}}]$ $=\mathbb{E}[\mathbb{E}[\xi_{n+1}|\mathcal{F}_{n}]M_{n-1}\mathbf{1}_{\Omega_{n}}]=0,$
@@
Using $(1-\gamma_{n}c)<1$ and putting all together we obtain the claim
@@
Lemma D.7.
Let $\delta>0$ be a tolerance level and $\gamma_{n}\leq\gamma_{1}$ be sufficiently small such that $\sum_{n=1}^{\infty}\gamma_{n}^{2}<\frac{\delta\epsilon}{2(G^{2}C^{2}+G^{2}+C)}$ and the condition in Lemma D.6 is fulfilled. Then, we have
@@
On the other hand it follows from Lemma D.10 that
@@
Lemma D.8.
Let $\delta>0$ be a tolerance level and $\gamma_{n}\leq\gamma_{1}$ be sufficiently small such that the condition in Lemma D.6 and Lemma D.7 are fulfilled. Moreover, we suppose $\gamma_{1}$ small enough such that $\frac{4C}{\mathbf{r}^{2}}\sum_{k=1}^{n}\gamma_{k}^{2}\leq\frac{\delta}{2}$. Then, we have
@@
By Lemma D.8, we have $\mathbb{P}(E_{n})\geq 1-\frac{\delta}{2}$. Moreover, by the additional step size assumption and Markov’s inequality we deduce that
@@
Hence, using Lemma D.8 we obtain
@@
(ii): We define $\tilde{D}_{n}:=D_{n}\mathbf{1}_{\Omega_{n}}$ and prove that $\tilde{D}_{n}\in o(1/n^{1-\eta})$, then the claim follows since $\mathbf{1}_{\Omega_{\mathcal{U}}}\leq\mathbf{1}_{\Omega_{n}}$ almost surely.
From the proof of Lemma D.4 and Lemma D.6 we have
$\tilde{D}_{n+1}\leq\tilde{D}_{n}-\gamma_{n}c\tilde{D}_{n}^{2\beta}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n}\rVert^{2}\mathbf{1}_{\Omega_{n}}.$
@@
$\leq\tilde{D}_{n}-\gamma_{n}c\tilde{D}_{n}^{2\beta}+\frac{LC}{2}\gamma_{n}^{2},$
where we have used that $D_{n}$ and $\mathbf{1}_{\Omega_{n}}$ are $\mathcal{F}_{n}$-measurable and $E[\lVert V_{n+1}(X_{n})\rVert^{2}|\mathcal{F}_{n}]\leq C$ from Equation ABC with $A=B=0$. By our step size choice we can apply Lemma 3.1 to obtain Claim (ii).
@@
$\beta=\frac{1}{2}$: We have from Lemma D.4 Equation 30 and Lemma D.6 that
$\tilde{D}_{n+1}$ $\leq(1-\gamma_{n}c^{2})\tilde{D}_{n}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}\mathbf{1}_{\Omega_{n}}.$
@@
where we used Equation 35 in the first inequality. By our choice of $\gamma_{n}$ there exists $\tilde{c}>0$ and $N>0$ such that $\gamma_{n}c^{2}-\frac{1-\eta}{n}+\frac{(1-\eta)\gamma_{n}c^{2}}{n}\geq\tilde{c}\gamma_{n}$ for all $n\geq N$. Thus, for all $n\geq N$
@@
by Equation 24 in Lemma 3.1 Hence, we apply Lemma A.3 to prove that $\lim_{n\to\infty}w_{n}=0$. By the definition of $w_{n}$ we have verified that $\mathbb{E}[(f(X_{n})-l)\mathbf{1}_{\Omega_{\mathcal{U}}}]\leq\mathbb{E}[\tilde{D}_{n}]\in o(\frac{1}{n^{1-\eta}})$
$\beta\in(\frac{1}{2},1]$: From Lemma D.4 Equation 31 and Lemma D.6 we have
$\tilde{D}_{n+1}$ $\leq(1-\gamma_{n}^{q}c^{2})\tilde{D}_{n}+\tilde{c}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+\gamma_{n}\xi_{n+1}\mathbf{1}_{\Omega_{n}}+\frac{L\gamma_{n}^{2}}{2}\lVert V_{n+1}(X_{n})\rVert^{2}\mathbf{1}_{\Omega_{n}},$
for $\tilde{c}=(2\beta)^{-\frac{1}{2\beta-1}}(1-\frac{1}{2\beta})c^{2}$. Next we multiply with $(n+1)^{1-\eta}$ and use Equation 17 to obtain
@@
$\leq(n+1)^{1-\eta}(1-\gamma_{n}^{q}c^{2})\mathbb{E}[\tilde{D}_{n}]+(n+1)^{1-\eta}\tilde{c}\gamma_{n}^{\frac{2\beta q-1}{2\beta-1}}+(n+1)^{1-\eta}\frac{LC}{2}\gamma_{n}^{2}$
@@
because we choose the auxiliary parameter $q$ as in the proof of Lemma 3.1 where we showed in Equation 24 and Equation 25 that
@@
The proof will be similar to the previous section. Instead of assuring that $(X_{n})$ remains close to the set where we could guarantee the unified gradient domination property, it is now sufficient that $f(X_{n})$ remains close to $f^{\ast}$ by the different definition of gradient domination definition in $f^{\ast}$. This will simplify the proof. Moreover, we may again assume w.l.o.g. the uniform second moment bounds, i.e. $A=B=0$, instead of the more general Equation ABC condition by the same argument as above but on the level sets.
@@
In the following, we fix the notation $D_{n}:=f(X_{n})-f^{\ast}$ and $\tilde{D}_{n}:=D_{n}\mathbf{1}_{\Omega_{n}}$, $\xi_{n+1}:=-\langle\nabla f(X_{n}),Z(X_{n},\zeta_{n+1})\rangle$ and obtain the parallel result to Lemma D.4.
Lemma D.9.